OpenStudy (anonymous):

Emily brought some money on a trip to the aquarium. She used half of the money to buy an admission ticket. Then, she used eight dollars to buy lunch at the food court. She used three-fourths of what was left to buy a souvenir, leaving her with four dollars. How much money did Emily take to the aquarium? forty-eight dollars forty dollars thirty-six dollars thirty-two dollars

OpenStudy (anonymous):

16*3/4= 12 and 16-12=4 dollars so (8+16)*2=48 dollars The Correct Answer is FORTY-EIGHT DOLLARS